# Why Chain of Thought Fails in Clinical Text Understanding

### v1

@@ 1 Introduction
Large language models (LLMs) are increasingly being explored for clinical workflows, including clinical decision support, documentation drafting, information extraction, and telemedicine communication (Park et al., 2024; Vrdoljak et al., 2025). However, deploying LLMs in healthcare, an inherently safety-critical setting, requires both high accuracy and clear interpretability (Habli et al., 2020; Ennab & Mcheick, 2024). Providing explicit and transparent reasoning allows clinicians to critically evaluate LLM decision-making processes, identify gaps or misapplications of evidence, and make more informed, trustworthy decisions in patient care (Amann et al., 2020). Therefore, reliability and transparency are crucial for developing trustworthy clinical AI tools and their integration into clinical practice (Kim et al., 2025).
Chain-of-thought (CoT) prompting (Wei et al., 2022), which encourages language models to generate intermediate reasoning steps prior to producing a final answer, has become a widely adopted technique for exposing intermediate rationales. Using explicit instruction, e.g., “Let’s think step by step”, CoT prompting can unlock the latent reasoning capabilities of LLMs, substantially improving performance and perceived interpretability across various domains. This effect is especially pronounced for tasks requiring advanced reasoning and logical coherence, such as mathematical problem-solving and coding (Kojima et al., 2022).
However, the effectiveness of CoT in clinical contexts remains underexplored and poorly understood. Existing evidence largely stems from exam-style or open-domain settings (e.g., USMLE-style QA) (Wu et al., 2024a; Liévin et al., 2024) rather than real-world, multi-source clinical narratives that demand heavy factual grounding and strict differential reasoning. The clinical text in electronic health records (EHRs) is noisy, ungrammatical, and highly variable, creating additional challenges for understanding. Moreover, previous studies have revealed that LLM outputs can contain hallucinations and incorrect information, potentially introducing patient-safety risks (Asgari et al., 2025; Roustan et al., 2025). Recent analyses further caution that CoT explanations may be unfaithful, representing superficially plausible narratives that do not reflect the model’s actual decision basis, inviting over-trust if used as a reliability signal (Lyu et al., 2023). These open questions motivate a dedicated, large-scale assessment of CoT clinical text understanding.
@@ 2.1 Large Language Models and Chain-of-Thought
The large-scale pre-training and extensive fine-tuning equip LLM with rich world knowledge and logical reasoning capabilities, enabling them to generalize to broad tasks (Raiaan et al., 2024). Wei et al. (2022) demonstrated that CoT prompting can guide models to generate intermediate reasoning steps and significantly improve performance on tasks spanning arithmetic, commonsense, and symbolic reasoning. Then, a simple “Let’s think step by step” instruction can elicit zero-shot reasoning Kojima et al. (2022), and self-consistency Wang et al. (2023) further boosts performance by aggregating diverse reasoning paths. These techniques have become a common practice for providing detailed explanations and improving performance. Subsequently, more advanced frameworks are built on CoT: ReAct interleaves reasoning with tool use (Yao et al., 2023b); Tree-of-Thoughts (Yao et al., 2023a) and Graph-of-Thoughts Besta et al. (2024) integrate structured planning and search with CoT; Least-to-Most (Zhou et al., 2023) decomposes complex problems into solvable subproblems. Furthermore, reinforcement-learning-based reasoning models (e.g., OpenAI o1 (Jaech et al., 2024) and DeepSeek-R1 (Guo et al., )) report additional gains by training models to “think before answering,” though these advances are demonstrated mainly on math, coding, or synthetic reasoning benchmarks rather than safety-critical clinical text.
@@ 2.2 Medical Reasoning
The clinical applications of LLMs have rapidly emerged. Notably, Med-PaLM-2 (Singhal et al., 2025), integrating base-model improvements with prompting/ensemble strategies inspired by CoT, reports expert-level performance across several medical examinations. Meanwhile, In-Context Padding (Wu et al., 2024b) infers important medical entities from the input and uses them to steer the model’s reasoning trajectory, improving alignment with clinical decision pathways. CLINICR (Nachane et al., 2024) mirrors the prospective process of incremental reasoning to support differential diagnosis. These methods aim to make intermediate steps clinically grounded rather than purely free-form rationales. However, Wu et al. (2024a) and Liévin et al. (2024) also revealed that CoT does not consistently improve the performance of general LLMs on medical questions. But there is still not a systematic, large-scale evaluation of CoT on real-world clinical text tasks beyond exam-style QA. Furthermore, safety-oriented assessments in clinical documentation and summarization report hallucinations (Shah, 2024), such as fabricated facts, incorrect dosages, misattributed conditions, and instruction-following failures with potential patient-safety implications, underscoring the need for domain-specific evaluation protocols that go beyond exam-style QA and measure factual grounding in real-world clinical text (Asgari et al., 2025).
@@ 3.2 Clinical Text Datasets
We include $87$ clinical text datasets from the BRIDGE benchmark (Wu et al., 2025). All tasks originate from real-world clinician–patient interactions drawn from electronic health records, clinical case reports, and online consultation records, with appropriate de-identification and license compliance.
@@ 3.4 Implementation Details
To examine LLMs under broadly adopted settings, the CoT prompt follows the classical CoT instruction (Wei et al., 2022; Wang et al., 2023; Singhal et al., 2025), instantiated as brief directives like “Solve it in a step by step fashion” and then state the final answer explicitly. We do not employ any prompt optimization to avoid introducing confounding factors that could obscure the core CoT effect.
Model access and inference setting.
@@ 4.3 Medical Concept Alignment Mitigates CoT Degradation
To examine how clinically grounded content relates to CoT reliability, we extract medical concepts from both the input $x_{i}$ and the CoT reasoning trace $r_{i}$ using cTAKES (Savova et al., 2010) (version 4.0.0.1), a clinical text analysis and knowledge extraction system that extracts clinical entities in free text and maps them into Concept Unique Identifiers (CUIs). For each instance, we define the set of input concepts $C_{x}$ and the set of concepts mentioned in the reasoning trace $C_{r}$ after normalization and de-duplication. We quantify alignment using two metrics to reflect the concept alignment:
@@
Figure 6: Lexical contrast between CoT-correct and CoT-incorrect traces. Each point is a word; axes show normalized frequency in two groups. Color encodes density from infrequent (yellow) to frequent (blue/red). The right panels list the most characteristic words for each group.
@@ 4.4 Lexical Signatures of CoT Reasoning Trace for Correct and Incorrect Instances
We partition the evaluation instances into two disjoint groups based on the correctness of the final answer under CoT: CoT-correct and CoT-incorrect. For each instance, we take the CoT reasoning trace $r_{i}$ and compute corpus-level word statistics after lowercasing, punctuation stripping, and lemmatization. To reduce confounding from input copying, we remove tokens that also appear in the corresponding input $x_{i}$. Let $V$ be the vocabulary after stopword removal. For a term $w\in V$, let $f_{w}^{\mathsf{cor}}$ and $f_{w}^{\mathsf{inc}}$ denote its token counts in CoT-correct and CoT-incorrect traces, with $N^{\mathsf{cor}}=\sum_{w}f_{w}^{\mathsf{cor}}$ and $N^{\mathsf{inc}}=\sum_{w}f_{w}^{\mathsf{inc}}$. We report (i) normalized frequencies $p_{w}^{\mathsf{cor}}=f_{w}^{\mathsf{cor}}/N^{\mathsf{cor}}$ and $p_{w}^{\mathsf{inc}}=f_{w}^{\mathsf{inc}}/N^{\mathsf{inc}}$ and (ii) differential usage via informative log-odds with a Dirichlet prior (Monroe et al., 2008; Kessler, 2017):
@@
where $\alpha_{w}$ is the background frequency scaled by $\alpha_{0}$ and $\sigma^{2}_{w}$ is the approximate variance from Monroe et al. (2008). This prior-regularized contrast mitigates small-sample artifacts and yields a ranked list of words characteristic of each group.
@@
Figure 6 contrasts term frequencies between CoT-correct and CoT-incorrect traces. Terms in the right-lower region (high in CoT-incorrect, low in CoT-correct) are disproportionately associated with errors. We observe that numbers, units, and measurement-related tokens are overrepresented among CoT-incorrect traces (e.g., pt, ct, numeric values), as are abbreviations for laboratory and physiologic assessments such as plt (platelet count) and pft (pulmonary function test). Conversely, CoT-correct traces contain a higher proportion of disease and therapy terms that are semantically aligned with the input clinical background. These lexical signatures align with our earlier analyses: tasks requiring precise handling of quantitative data and shorthand notations are more vulnerable under CoT, consistent with reports that LLMs are brittle with numeric reasoning (Akhtar et al., 2023; Khandekar et al., 2024). and domain-specific abbreviations (Rajkomar et al., 2022; Kugic et al., 2024).
@@ 4.5 Error Analysis: Hallucination, Omission, and Incompleteness
To characterize failure modes in the CoT reasoning trace $r$, we adopt an LLM-as-a-Judge protocol. Following prior studies about LLM-as-a-Judge on clinical task (Liu et al., 2024; Gu et al., 2024), we evaluate three primary dimensions:
@@
We adopt OpenAI-o3 as the judge model. The judge model was manually assessed by a clinical expert in a random subset ($n=200$). The full prompts can be found in Appendix A.5. To minimize bias, model identities are anonymized and responses are randomly ordered during judging. We analyze three challenging EHR-based tasks—mortality prediction on MIMIC-III-Outcome (van Aken et al., 2021), disease diagnosis on MIMIC-IV-DiReCT dataset (Wang et al., 2024) and MIMIC-IV-CDM dataset (Hager et al., 2024).
@@
Figure 7 reports the distribution of detected error types across models. Generally, the best-performing models, GPT-4o, show the fewest errors; the following Qwen2.5-72B-Instruct and Llama-3.3-70B-Instruct are comparable. However, MeLLaMA-70B-Chat, trained primarily for medical Q&A rather than EHR cases, exhibits substantially higher omission and incompleteness rates, indicating difficulty in extracting and organizing patient-specific evidence from longitudinal notes. Overall, hallucination and omission are the predominant failure modes (Figure 7). The expert review highlights cases where historical conditions mentioned in the input are incorrectly treated as current. Compared to expert annotation, the positive predictive value (PPV) of the judge was $97.22\%$ for hallucination, $97.83\%$ for omission, and $100\%$ for incompleteness, demonstrating the reliability of the LLM-as-a-Judge.

### v2

@@ 1 Introduction
Large language models (LLMs) are increasingly being explored for clinical workflows, including clinical decision support, documentation drafting, information extraction, and telemedicine communication (park2024assessing; vrdoljak2025review). However, deploying LLMs in healthcare, an inherently safety-critical setting, requires both high accuracy and clear interpretability (habli2020artificial; ennab2024enhancing). Providing explicit and transparent reasoning allows clinicians to critically evaluate LLM decision-making processes, identify gaps or misapplications of evidence, and make more informed, trustworthy decisions in patient care (amann2020explainability). Therefore, reliability and transparency are crucial for developing trustworthy clinical AI tools and their integration into clinical practice (kim2025transparency).
Chain-of-thought (CoT) prompting (cot-first-wei2022), which encourages language models to generate intermediate reasoning steps prior to producing a final answer, has become a widely adopted technique for exposing intermediate rationales. Using explicit instruction, e.g., “Let’s think step by step”, CoT prompting can unlock the latent reasoning capabilities of LLMs, substantially improving performance and perceived interpretability across various domains. This effect is especially pronounced for tasks requiring advanced reasoning and logical coherence, such as mathematical problem-solving and coding (cot-step-kojima2022).
However, the effectiveness of CoT in clinical contexts remains underexplored and poorly understood. Existing evidence largely stems from exam-style or open-domain settings (e.g., USMLE-style QA) (llm-cnmle; medical-reason-llm) rather than real-world, multi-source clinical narratives that demand heavy factual grounding and strict differential reasoning. The clinical text in electronic health records (EHRs) is noisy, ungrammatical, and highly variable, creating additional challenges for understanding. Moreover, previous studies have revealed that LLM outputs can contain hallucinations and incorrect information, potentially introducing patient-safety risks (asgari2025framework; roustan2025clinicians). Recent analyses further caution that CoT explanations may be unfaithful, representing superficially plausible narratives that do not reflect the model’s actual decision basis, inviting over-trust if used as a reliability signal (turpin2023faithfulness). These open questions motivate a dedicated, large-scale assessment of CoT clinical text understanding.
@@ 2.1 Large Language Models and Chain-of-Thought
The large-scale pre-training and extensive fine-tuning equip LLM with rich world knowledge and logical reasoning capabilities, enabling them to generalize to broad tasks (llm-review-2024). cot-first-wei2022 demonstrated that CoT prompting can guide models to generate intermediate reasoning steps and significantly improve performance on tasks spanning arithmetic, commonsense, and symbolic reasoning. Then, a simple “Let’s think step by step” instruction can elicit zero-shot reasoning cot-step-kojima2022, and self-consistency cot-self-consistency-wang2022 further boosts performance by aggregating diverse reasoning paths. These techniques have become a common practice for providing detailed explanations and improving performance. Subsequently, more advanced frameworks are built on CoT: ReAct interleaves reasoning with tool use (cot-react-yao2022); Tree-of-Thoughts (cot-tot-2023) and Graph-of-Thoughts cot-got-2023 integrate structured planning and search with CoT; Least-to-Most (cot-least-to-most-zhou2022) decomposes complex problems into solvable subproblems. Furthermore, reinforcement-learning-based reasoning models (e.g., OpenAI o1 (openai-o1) and DeepSeek-R1 (deepseek-r1)) report additional gains by training models to “think before answering,” though these advances are demonstrated mainly on math, coding, or synthetic reasoning benchmarks rather than safety-critical clinical text.
@@ 2.2 Medical Reasoning
The clinical applications of LLMs have rapidly emerged. Notably, Med-PaLM-2 (medical-medpalm2), integrating base-model improvements with prompting/ensemble strategies inspired by CoT, reports expert-level performance across several medical examinations. Meanwhile, In-Context Padding (icp-ijcai-2024) infers important medical entities from the input and uses them to steer the model’s reasoning trajectory, improving alignment with clinical decision pathways. CLINICR (llm-cot-medical-2024) mirrors the prospective process of incremental reasoning to support differential diagnosis. These methods aim to make intermediate steps clinically grounded rather than purely free-form rationales. However, llm-cnmle and medical-reason-llm also revealed that CoT does not consistently improve the performance of general LLMs on medical questions. Especially, cot-fail-ood demonstrated that the benefit of CoT reasoning vanishes when it is pushed beyond training distributions. However, there is still no systematic, large-scale evaluation of CoT on clinical text tasks. Furthermore, safety-oriented assessments in clinical documentation and summarization report hallucinations (llm-hallucination-ehr), such as fabricated facts, misattributed conditions, and instruction-following failures with potential patient-safety implications, underscoring the need for domain-specific evaluation protocols that go beyond exam-style QA and measure factual grounding in real-world clinical text (asgari2025framework).
@@ 3.2 Clinical Text Datasets
We include $87$ clinical text datasets from the BRIDGE benchmark (nejm-dataset-review; BRIDGE). All tasks originate from real-world clinician–patient interactions drawn from electronic health records, clinical case reports, and online consultation records, with appropriate de-identification and license compliance.
@@ 3.4 Implementation Details
To examine LLMs under broadly adopted settings, the CoT prompt follows the classical CoT instruction (cot-first-wei2022; cot-self-consistency-wang2022; medical-medpalm2), instantiated as brief directives like “Solve it in a step by step fashion” and then state the final answer explicitly. We do not employ any prompt optimization to avoid introducing confounding factors that could obscure the core CoT effect.
Model and inference setting.
@@ 4.1 CoT Undermines LLM Performance in Clinical Text Tasks
We further evaluate CoT under more favorable inference strategies. First, combining CoT with few-shot prompting across all 87 tasks and 7 representative models (Appendix A.4) confirms that few-shot examples substantially improve performance, but few-shot with CoT still does not yield additional gains (see S3 and S1). Second, integrating CoT with self-consistency on 27 classification tasks for 5 models (Appendix A.5) shows that ensembling multiple samples can boost model performance and narrow the gap between CoT and standard prompt without analysis, yet CoT remains generally weaker than standard prompt and incurs higher computational cost (see S4 and S2). Together, these results indicate that stronger prompting and decoding strategies can partially mitigate CoT-induced degradation, but the CoT still fails to improve interpretability without sacrificing performance.
Additionally, we also perform subgroup analyses across task types (Appendix A.6 and languages (Appendix A.7). Across eight task categories, CoT-induced degradation is widespread, with particularly large drops in event extraction, summarization, and question answering, while normalization/coding emerges as the only category with consistent (but modest) gains. Across nine languages, CoT also tends to reduce performance in most settings, indicating that the instability of vanilla CoT is a multilingual phenomenon rather than an English-only artifact, although the magnitude and even direction of the effect vary by language and model. These subgroup results reinforce our main conclusion that CoT is not reliably beneficial for real-world clinical text tasks, while highlighting specific niches (e.g., normalization/coding and certain language–model pairs) where step-by-step reasoning may offer limited benefits.
@@ 4.3 Medical Concept Alignment Mitigates CoT Degradation
To examine how clinically grounded content relates to CoT reliability, we extract medical concepts from both the input $x_{i}$ and the CoT reasoning trace $r_{i}$ using cTAKES (cTAKES) (version 4.0.0.1), a clinical text analysis and knowledge extraction system that extracts clinical entities in free text and maps them into Concept Unique Identifiers (CUIs). For each instance, we define the set of input concepts $C_{x}$ and the set of concepts mentioned in the reasoning trace $C_{r}$ after normalization and de-duplication. We quantify alignment using two metrics to reflect the concept alignment:
@@
As exemplified by our case study in Appendix A.8, CoT traces produced by high-capacity models hallucinate or distort key clinical facts (e.g., misinterpreting some numerical clinical features), driven by misunderstandings of clinical abbreviations and numerical values. These errors reflect weak grounding in the numerical features and medical concepts frequently contained in clinical text.
Figure 6: Lexical contrast between correct and incorrect CoT responses. Each point corresponds to a word, with axes showing its normalized frequency in the correct vs. incorrect groups. Color encodes relative frequency: blue indicates higher frequency in correct CoT, red indicates higher frequency in incorrect CoT, and yellow indicates similar frequency in both groups. The right panels list the most characteristic words for each group.
@@ 4.4 Lexical Signatures of CoT Reasoning Trace for Correct and Incorrect Instances
We partition the evaluation instances into two disjoint groups based on the correctness of the final answer under CoT: CoT-correct and CoT-incorrect. For each instance, we take the CoT reasoning trace $r_{i}$ and compute corpus-level word statistics after lowercasing, punctuation stripping, and lemmatization. To reduce confounding from input copying, we remove tokens that also appear in the corresponding input $x_{i}$. Let $V$ be the vocabulary after stopword removal. For a term $w\in V$, let $f_{w}^{\mathsf{cor}}$ and $f_{w}^{\mathsf{inc}}$ denote its token counts in CoT-correct and CoT-incorrect traces, with $N^{\mathsf{cor}}=\sum_{w}f_{w}^{\mathsf{cor}}$ and $N^{\mathsf{inc}}=\sum_{w}f_{w}^{\mathsf{inc}}$. We report (i) normalized frequencies $p_{w}^{\mathsf{cor}}=f_{w}^{\mathsf{cor}}/N^{\mathsf{cor}}$ and $p_{w}^{\mathsf{inc}}=f_{w}^{\mathsf{inc}}/N^{\mathsf{inc}}$ and (ii) differential usage via informative log-odds with a Dirichlet prior (word-diff-principle; word-diff-scatterplot):
@@
where $\alpha_{w}$ is the background frequency scaled by $\alpha_{0}$ and $\sigma^{2}_{w}$ is the approximate variance from word-diff-principle. This prior-regularized contrast mitigates small-sample artifacts and yields a ranked list of words characteristic of each group.
@@
Figure 6 contrasts term frequencies between CoT-correct and CoT-incorrect traces. Terms in the right-lower region (high in CoT-incorrect, low in CoT-correct) are disproportionately associated with errors. We observe that numbers, units, and measurement-related tokens are overrepresented among CoT-incorrect traces (e.g., pt, ct, numeric values), as are abbreviations for laboratory and physiologic assessments such as plt (platelet count) and pft (pulmonary function test). Conversely, CoT-correct traces contain a higher proportion of disease and therapy terms that are semantically aligned with the input clinical background. These lexical signatures align with our earlier analyses: tasks requiring precise handling of quantitative data and shorthand notations are more vulnerable under CoT, consistent with reports that LLMs are brittle with numeric reasoning (llm-numerical-reasoning-2023; llm-numerical-reasoning-med). and domain-specific abbreviations (llm-abbr-medical-nc; llm-abbr-medical-jamia).
@@ 4.5 Error Analysis: Hallucination, Omission, and Incompleteness
To characterize failure modes in the CoT reasoning trace $r$, we adopt an LLM-as-a-Judge protocol. Following prior studies about LLM-as-a-Judge on clinical task (chatgpt-eye-diagnosis; llm-judge-review), we evaluate three primary dimensions:
@@
We adopt OpenAI-o3 as the judge model. The judge model was manually assessed by a clinical expert in a random subset ($n=200$). The full prompts can be found in Appendix A.10. To minimize bias, model identities are anonymized and responses are randomly ordered during judging. We analyze three challenging EHR-based tasks—mortality prediction on MIMIC-III-Outcome (dataset-MIMIC-III-Outcome), disease diagnosis on MIMIC-IV-DiReCT dataset (dataset-MIMIC-IV-DiReCT) and MIMIC-IV-CDM dataset (dataset-MIMIC-IV-CDM).
@@
Figure 7 reports the distribution of detected error types across models. Generally, the best-performing models, GPT-4o, show the fewest errors; the following Qwen2.5-72B-Instruct and Llama-3.3-70B-Instruct are comparable. However, MeLLaMA-70B-Chat, trained primarily for medical Q&A rather than EHR cases, exhibits substantially higher omission and incompleteness rates, indicating difficulty in extracting and organizing patient-specific evidence from longitudinal notes. Overall, hallucination and omission are the predominant failure modes (Figure 7). The expert review highlights cases where historical conditions mentioned in the input are incorrectly treated as current. Compared to expert annotation, the positive predictive value (PPV) of the judge was $97.22\%$ for hallucination, $97.83\%$ for omission, and $100\%$ for incompleteness, demonstrating the reliability of the LLM-as-a-Judge. As the case study (Appendix A.8) illustrates that these patterns in concrete EHR scenarios, where CoT traces hallucinate or distort key clinical details and omit salient findings, align well with the identified hallucination, omission, and incompleteness.